OpenStudy (anonymous):

Solve: 3(x + 4) greater than or equal to x - 2 3x + 12 greater than or equal to x - 2 what do i do next?

OpenStudy (anonymous):

\[3x + 12 \ge x - 2 \]

OpenStudy (anonymous):

get rid of the x on the right side by subtracting x from both sides

OpenStudy (anonymous):

\[2x + 12 \ge -2 \]

OpenStudy (anonymous):

subtract 12 from both sides

OpenStudy (anonymous):

you mean : \(3(x+4) \ge x-2 \) and : \(3x+12\ge x-2\) right ?

OpenStudy (anonymous):

subtract 12 from both sides

OpenStudy (anonymous):

\[2x \ge -14\]

OpenStudy (anonymous):

divide both sides by 2

OpenStudy (anonymous):

\[x \ge -7\]

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

yep